# A comparison between geostatistical and machine learning models for spatio-temporal prediction of PM2.5 data

### v1

@@ Abstract
Ambient air pollution poses significant health and environmental challenges. Exposure to high concentrations of PM2.5 have been linked to increased respiratory and cardiovascular hospital admissions, more emergency department visits and deaths. Traditional air quality monitoring systems such as EPA-certified stations provide limited spatial and temporal data. The advent of low-cost sensors has dramatically improved the granularity of air quality data, enabling real-time, high-resolution monitoring. This study exploits the extensive data from PurpleAir sensors to assess and compare the effectiveness of various statistical and machine learning models in producing accurate hourly PM2.5 maps across California. We evaluate traditional geostatistical methods, including kriging and land use regression, against advanced machine learning approaches such as neural networks, random forests, and support vector machines, as well as ensemble model. Our findings enhanced the predictive accuracy of PM2.5 concentration by correcting the bias in PurpleAir data with an ensemble model, which incorporating both spatiotemporal dependencies and machine learning models.
@@
Keywords: Air pollution; Geostatistical model; Spatio-temporal data; Machine learning model; Ensemble model
@@ 1 Introduction
The disparate distribution of air pollution underscores significant ecological justice concerns Hajat et al. (2015), particularly with particulate matter less than 2.5 micrometers (PM2.5). As one of the most hazardous pollutants, it not only leads to significant health risks but also contributes to environmental issues Organization et al. (2018). In urban settings, the government typically installs a few EPA-certified air quality monitors that are updated hourly with an hourly average particle count. However, pollution levels exhibit considerable variability within small areas and over short periods. The advent of low-cost air pollution sensors, such as PurpleAir, has revolutionized monitoring by providing high spatiotemporal resolution data, thereby enhancing the detection of pollution hotspots and improving air quality indices during severe pollution events like wildfires Morawska et al. (2018); Bi et al. (2020); Delp and Singer (2020).
@@
PurpleAir sensors, known for their affordability, ease of use and maintenance, have proliferated, significantly aiding public and scientific understanding of air quality dynamics Barkjohn et al. (2020). The PurpleAir sensor network offers a high spatial resolution and real-time particulate matter data at a 2-minute temporal resolution. The widespread use of PurpleAir sensors, gives us the opportunity to generate an hourly map of ambient PM2.5 concentrations in California. Leveraging high spatiotemporal resolution data collected from 1110 PurpleAir sensors operating from January to December 2019, this study aims to assess various geostatistical and machine learning techniques in producing hourly PM2.5 maps.
@@
In recent years, considerable effort has been devoted to developing statistical methods that accurately predict PM2.5 concentrations on the spatial, temporal and spatiotemporal domains. Traditional statistical methods like kriging and land use regression have been essential for epidemiological studies to estimate PM2.5{2.5} levels. However, these methods often suffer from computational intensity and data availability limitations, respectively Alexeeff et al. (2015); Datta et al. (2016a); Hu et al. (2013). With the increase in data volume, computational challenges have prompted the adoption of machine learning models, which have demonstrated superior performance in capturing complex patterns in high-dimensional data. Datta et al. (2016a) introduced a hierarchical nearest-neighbor Gaussian process model for large geostatistical datasets, which used a hierarchical structure to reduce the computational complexity of the Gaussian process while maintaining its accuracy. However, it requires spatiotemporal covariance structures and parametric assumptions.
More recently, machine learning (ML) algorithms have been explored for air pollution prediction, including random forests, support vector regression, and deep learning approaches. Hu et al. (2017) combined ground observations, satellite data, and meteorological variables to develop a model that could predict PM2.5 concentrations at a high spatial resolution using a random forest (RF). The study was able to accurately estimate PM2.5 concentrations across the United States, though they found that season and climate can affect the prediction accuracy. Mogollón-Sotelo et al. (2021) adopted support vector machines (SVM) to forecast PM2.5 concentrations in Bogotá, Colombia. The study trained data for 12 monitoring stations using a radial basis kernel, and the model had better predictive accuracy compared with Bogotá Integrated Air Quality Modeling System, especially in areas with complex terrain, where traditional air quality models did not perform well. Gupta and Christopher (2009) used a neural network to produce more accurate predictions of PM2.5 concentrations compared with other traditional regression models. The method involved training a neural network to predict PM2.5 concentrations based on the input of three years of surface, satellite, and meteorological fields in the southeastern United States, resulting in significant improvements in accuracy compared to traditional regression-based methods.
@@
Despite the variety of approaches available, there is no consensus on which produces the most accurate predictions. Several studies have evaluated traditional geostatistical methods with machine learning algorithms. Requia et al. (2019) compared ordinary kriging, hybrid interpolation, and random forests for estimating concentrations of 10 PM2.5 components. And domenstrated that random forests surpassed both kriging and hybrid interpolation methods (empirical Bayesian kriging and land use regression).Wang et al. (2019) introduced a flexible spatial prediction approach using a Nearest-Neighbor Neural Network, which excelled over conventional geostatistical methods in handling non-normal data by innovatively utilizing neighboring information. Conversely, Berrocal et al. (2020) evaluated universal kriging and downscaler models against machine learning approaches for generating daily national maps of PM2.5 concentrations in the United States, noting superior predictive performance in geostatistical methods under the assumption of constant spatial covariance.
@@
Although numerous studies have compared geostatistical and machine learning techniques, they often focus solely on spatial or temporal dependencies, neglecting their combined spatiotemporal dynamics. Leveraging the extensive network of PurpleAir sensors, this study aims to produce hourly maps for PM2.5 concentrations in California. We explore the efficacy of both traditional geostatistical methods, such as universal kriging and fixed rank kriging, and advanced non-geostatistical approaches, including regression, random forests, support vector machines, and neural networks. Additionally, this study integrates geostatistical methods and nearest neighbor data with machine learning algorithms to enhance prediction accuracy and performance.
@@ 2 Data
PurpleAir sensors are one of the low-cost particulate matter (PM) sensors that have gained popularity for their affordability, accessibility, and ease of use in monitoring outdoor air quality. These sensors use laser-based optical sensors to count particles by sizes and uses the counts to calculate mass concentrations of PM1.0, PM2.5, and PM10. Each sensor has two Plantower PMS5003 unites,, labeled as channels A and B, which operate alternately and provide 80-second averaged data values (and, from June 2019, 2-minute averaged vales) (Barkjohn et al., 2020). Widely adopted by communities, citizen scientists, and researchers, PurpleAir sensors are extensively used for local air quality monitoring, and their data has been utilized in various applications. The substantial rise in sensor counts from 251 in 2017 to 39266 in 2022 (API, ) as shown in Figure 1, reflects a growing interest in understanding air quality at a local level. This increase has important implications for air quality monitoring and decision-making. Figure 2 illustrates the rapid increase in PurpleAir installations across California from 2019 to 2023. This underscores the growing importance of incorporating these sensors into air quality monitoring efforts. With more sensors in operation, a higher density of data points can be obtained, enabling a more comprehensive and detailed assessment of air quality trends and patterns in specific areas.
Figure 1: PurpleAir sensors counts from 2017-2019 in US
Figure 2: Real-time map for PM2.5 concentration measured by PurpleAir sensors in 2019 (left) versus 2023 (right) PurpleAir (2021). (Green) indicates air quality is satisfactory, (yellow) air quality is acceptable, (orange) members from sensitive maybe affected, (red) the general public may experience health effects, and (purple) risk of health effects is increased for everyone.
However, it is essential to consider the limitations of PurpleAir sensors, such as potential issues with sensor accuracy, calibration, and data quality. As PurpleAir sensor counts continue to rise, it is crucial to ensure that proper quality control measures, calibration, and validation protocols are in place to ensure the reliability and accuracy of the data obtained from these sensors for robust air quality assessments and decision-making. To start with the most robust data we processed the PurpleAir data in three steps after accessing and downloading the data in JSON format: 1) we removed observations with extremely high or low temperatures (i.e 2147483447 or -224) which indicates a communication error between PurpleAir micro-controller and temperature sensors. We also removed observations that have a relative humidity (RH) outside the range $0-100\%$. 2) We checked the consistency of the measurements for both channels A and B and then averaged the 2-minute or 80-second data hourly. 3) we applied the correction formula given by Barkjohn et al. (2020),
@@
where PM2.5corrected is the corrected PM2.5, $PA_{avg}$ is the hourly averaged PM2.5 from channel A and B and $RH$ is hourly relative humidity. For our study, We collected hourly measurements of PM2.5 concentrations from January 2019 through to December 31, 2019, from 1015 PurpleAir sensors in California. Figure 3 illustrates an example of the data.
@@ 3 Methods
3.1 Geostatisticsl Methods
Geostatistical methods are a set of statistical techniques that are used to analyze spatial/spatiotemporal data, explicitly take into account the spatial/spatiotemporal dependence of the data. These methods are widely used in various fields such as geology, hydrology, agriculture, environmental science, and many others (Maliva, 2016). A typical procedure in geostatistical techniques is to start with modeling the spatial/spatiotemporal correlation structure of the data, which describes how the values at different locations are related to each other in space and over time. Once we estimate the model parameter from the data, we can make predictions for the variable of interest at unsampled locations.
3.1.1 Universal Kriging (UK)
@@
where, $\mu(s,t)=\mbox{\bf X}(s,t)\mbox{\boldmath$\beta$}$ is the mean with $\mbox{\bf X}(s,t)=[X_{1}(s,t),X_{2}(s,t),\dots,X_{p}(s,t)]^{T}$ are predictors and $\beta$ is the $p$-dimensional vector of parameters. $w(s,t)$ represents a mean-zero Gaussian process with a spatiotemporal covariance function. $\epsilon(s,t)$ is the white noise measurement error with mean 0 and variance $\sigma_{\epsilon}^{2}$. We use a separable exponential covariance function,
@@
where $\mbox{\bf h}=s^{\prime}-s$ is the spatial lag, $\tau=t^{\prime}-t$ temporal lag and $\mbox{\boldmath$\theta$}=\{\sigma_{s},\sigma_{t},\rho_{s},\rho_{t}\}$ is the set of parameters associated with this covariance function, where $\sigma_{s},\sigma_{t}$ are the standard deviations for space and time , and $\rho_{s},\rho_{t}$ are spatial and time range parameters respectively. To predict Y at any unsampled location at $(s_{0},t_{0})$, we have
$\hat{Y}(s_{0},t_{0})=\mbox{\bf X}(s_{0},t_{0})^{T}\hat{\mbox{\boldmath$\beta$}}_{gls}+c(\mbox{\boldmath$\theta$})^{T}\Sigma_{y}^{-1}(\mbox{\boldmath$\theta$})(\mbox{\bf Y}-\mbox{\bf X}\hat{\mbox{\boldmath$\beta$}}_{gls}),$ (3)
where $c(\mbox{\boldmath$\theta$})=\text{Cov}(Y(s_{0},t_{0}),\mbox{\bf Y})$ and $\text{Var}(\mbox{\bf Y})=\Sigma_{y}(\mbox{\boldmath$\theta$})$. The uncertainty can be obtained by
$\text{Var}(\hat{Y}(s_{0},t_{0}))=c_{0,0}-c(\mbox{\boldmath$\theta$})^{T}\Sigma_{y}^{-1}(\mbox{\boldmath$\theta$})c(\mbox{\boldmath$\theta$})+\kappa,$ (4)
@@
$\kappa=\left(\mbox{\bf X}(s_{0},t_{0})-\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{\boldmath$\theta$})c(\mbox{\boldmath$\theta$})\right)^{T}\left(\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{\boldmath$\theta$})\mbox{\bf X}\right)^{-1}\left(\mbox{\bf X}(s_{0},t_{0})-\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{\boldmath$\theta$})c(\mbox{\boldmath$\theta$})\right)$.
The above prediction method is called universal kriging. To get the prediction in (3) and prediction uncertainty in (4), we have to compute the estimates $\hat{\mbox{\boldmath$\beta$}}_{gls}$ and $\hat{\mbox{\boldmath$\theta$}}$ so that we get the estimated BLUP (EBLUP). A 95% prediction coverage is given by
$\hat{Y}(s_{0},t_{0})\pm 1.96\times\sqrt{\text{Var}(\hat{Y}(s_{0},t_{0}))}.$ (5)
In this study, we implemented universal kriging using the gstat package in R Gräler et al. (2016).
3.1.2 Nearest Neighbor Gaussian Process (NNGP)

### v2

@@ Abstract
Ambient air pollution poses significant health and environmental challenges. Exposure to high concentrations of PM2.5 has been linked to increased respiratory and cardiovascular hospital admissions, more emergency department visits and deaths. Traditional air quality monitoring systems such as EPA-certified stations provide limited spatial and temporal data. The advent of low-cost sensors has dramatically improved the granularity of air quality data, enabling real-time, high-resolution monitoring. This study exploits the extensive data from PurpleAir sensors to assess and compare the effectiveness of various statistical and machine learning models in producing accurate hourly PM2.5 maps across California. We evaluate traditional geostatistical methods, including universal kriging, nearest neighbor Gaussian process and fixed rank kriging, against advanced machine learning approaches such as neural network, random forest, and support vector machine, as well as ensemble model. Our findings highlight the synergistic value of combining geostatistical methods with machine learning to achieve higher-accuracy PM2.5 predictions with lower computational burden, while flexibly capturing nonlinear structure when present.
@@
Keywords: Air pollution; Geostatistical model; Spatio-temporal data; Machine learning model; Ensemble model f
@@ 1 Introduction
The disparate distribution of air pollution underscores significant ecological justice concerns (Hajat et al., 2015), particularly for fine particulate matter with an aerodynamic diameter less than 2.5 micrometers (PM2.5). As one of the most hazardous air pollutants, PM2.5{2.5} is associated with substantial adverse health effects and contributes to broader environmental risks (World Health Organization, 2018). In urban settings, the government typically operates a limited number of United States Environmental Protection Agency (EPA)–certified air quality monitoring stations, which report hourly averaged pollutant concentrations. However, air pollution levels can vary substantially over short temporal scales and within small spatial areas. The emergence of low-cost air pollution sensors, such as PurpleAir (PA) sensors, has substantially expanded monitoring coverage by providing measurements at high spatial and temporal resolution. These data have improved the identification of localized pollution hotspots and enhanced the assessment of air quality conditions often summarized using the Air Quality Index (AQI) during extreme pollution events such as wildfires (Morawska et al., 2018; Bi et al., 2020; Delp and Singer, 2020).
@@
PurpleAir data offer dense spatial coverage and minute-level sampling, enabling hourly mapping of ambient PM2.5 across California with substantially finer spatial granularity than the EPA reference network (Barkjohn et al., 2020). Leveraging high spatiotemporal resolution data collected from 1015 PurpleAir sensors operating from January to December 2019, this study aims to assess various geostatistical and machine learning techniques in producing hourly PM2.5 maps.
@@
In recent years, considerable effort has been devoted to developing statistical methods that accurately predict PM2.5 concentrations on the spatial, temporal, and spatiotemporal domains. Traditional statistical methods like kriging and land use regression have been essential for epidemiological studies to estimate PM2.5{2.5} levels. However, these methods often suffer from computational intensity and data availability limitations, respectively (Alexeeff et al., 2015; Hu et al., 2013). Datta et al. (2016a) introduced a hierarchical nearest-neighbor Gaussian process model for large geostatistical datasets, which used a hierarchical structure to reduce the computational complexity of the Gaussian process while maintaining its accuracy. However, it requires spatiotemporal covariance structures and parametric assumptions. Beyond classical frameworks, Bayesian geostatistical models implemented via integrated nested Laplace approximations (INLA) with stochastic partial differential equations (SPDEs) representations provide scalable inference on Gaussian random fields and full predictive uncertainty, and have been applied to PM2.5 mapping in recent work (Gong et al., 2021; Chen et al., 2023). Implementing INLA would require a different modeling pipeline (SPDE mesh, priors), so we position it as state-of-the-art context rather than a head-to-head comparator in this study.
More recently, machine learning algorithms have been explored for air pollution prediction, including random forest, support vector regression, and deep learning approaches. Representative studies combine ground observations, satellite products, and meteorology to improve accuracy at high spatial resolution (e.g., random forest, support vector regression and neural network frameworks), while geostatistical baselines remain competitive under appropriate covariance assumptions (Hu et al., 2017; Mogollón-Sotelo et al., 2021; Gupta and Christopher, 2009; Berrocal et al., 2020). Taken together, the current literature spans two paradigms: probabilistic geostatistical models that explicitly encode spatial dependence and provide predictive uncertainty, and flexible machine learning models that capture nonlinearity but typically yield point predictions.
@@
Despite the variety of approaches available, there is no consensus on which produces the most accurate predictions. Several studies have evaluated traditional geostatistical methods with machine learning algorithms. Requia et al. (2019) compared ordinary kriging, hybrid interpolation, and random forest for estimating concentrations of 10 PM2.5 components. They demonstrated that random forest surpassed both kriging and hybrid interpolation methods (empirical Bayesian kriging and land use regression). Wang et al. (2019) introduced a flexible spatial prediction approach using a nearest neighbor neural network, which excelled over conventional geostatistical methods in handling non-normal data by innovatively utilizing neighboring information. Conversely, Berrocal et al. (2020) evaluated universal kriging and downscaler models against machine learning approaches for generating daily national maps of PM2.5 concentrations in the United States, noting superior predictive performance in geostatistical methods under the assumption of constant spatial covariance.
@@
Although numerous studies have compared geostatistical and machine learning techniques, they often focus solely on spatial or temporal dependencies, neglecting their combined spatiotemporal dynamics. Leveraging the extensive network of PurpleAir sensors, this study aims to produce hourly maps for PM2.5 concentrations in California. We explore the efficacy of both traditional geostatistical methods, such as universal kriging, nearest neighbor Gaussian process and fixed rank kriging, and advanced non-geostatistical approaches, including regression, random forest, support vector machines, and neural networks. Additionally, this study integrates geostatistical methods and nearest neighbor data with machine learning algorithms to enhance prediction accuracy and performance. Methodologically, we evaluate short-range interpolation performance in a dense-sensor setting and quantify gains from hybridizing kriging-based predictors with machine learning models. Practically, the resulting hourly maps and associated uncertainty where available can support regulatory and operational tasks such as hotspot surveillance during wildfire events, AQI gap-filling in areas without reference monitors. The remainder of the article is organized as follows. Section 2 describes the motivating dataset. Section 3 outlines the methodological framework. Section 4 presents the data analysis and results. Section 5 concludes with a summary of the findings, limitation of the current work and directions for future work.
@@ 2 Data
PurpleAir sensors are one of the low-cost particulate matter (PM) sensors that have gained popularity for their affordability, accessibility, and ease of use in monitoring outdoor air quality. These sensors use laser-based optical sensors to count particles by sizes and use the counts to calculate the mass concentrations of PM1.0, PM2.5, and PM10. Each sensor has two Plantower PMS5003 units, labeled as channels A and B, which operate alternately and provide 80-second averaged data values (and, from June 2019, 2-minute averaged values) (Barkjohn et al., 2020). Beyond mass, PurpleAir also records meteorological covariates (temperature in °F, relative humidity (RH) in %, and barometric pressure in hPa), timestamps, and geolocation (latitude/longitude). Widely adopted by communities, citizen scientists, and researchers, PurpleAir sensors are extensively used for local air quality monitoring, and their data has been utilized in various applications. The substantial increase in sensor counts from 251 in 2017 to 39266 in 2022 (1) as shown in Figure 1, reflects a growing interest in understanding air quality at the local level. This increase has important implications for air quality monitoring and decision-making. Figure 2 illustrates the rapid increase in PurpleAir installations in California from 2019 to 2023. This underscores the growing importance of incorporating these sensors into air quality monitoring efforts. With more sensors in operation, a higher density of data points can be obtained, allowing a more comprehensive and detailed assessment of air quality trends and patterns in specific areas.
Figure 1: PurpleAir sensors counts from 2017-2022 in US
Figure 2: Real-time map of PM2.5 concentration measured by PurpleAir sensors in 2019 (left) and 2023 (right), reproduced from the PurpleAir public visualization interface (PurpleAir, 2021). Marker colors indicate air quality categories (green) satisfactory; (yellow) acceptable; (orange) members of sensitive groups may be affected; (red) the general public may experience health effects; (purple) increased health risk for everyone. Marker sizes reflect visualization rendering and display resolution in the original interface and do not represent the number or density of installed sensors.
[…]
@@
where PM2.5corrected is the corrected PM2.5, $PA_{avg}$ is the hourly averaged PM2.5 from channel A and B and $RH$ is hourly relative humidity. For our study, we collected hourly measurements of PM2.5 concentrations from January 2019 through to December 31, 2019, from 1015 PurpleAir sensors in California, of which 1013 remained after QC and were used in modeling. Figure 3 illustrates an example of the data.
@@ 3 Methods
3.1 Overview
Our methodological framework combines three geostatistical models, universal kriging (UK), fixed rank kriging (FRK), and nearest neighbor Gaussian process (NNGP), and several non-geostatistical machine learning models, including random forest (RF), support vector regression(SVR), and neural network (NN). These models were selected to represent distinct theoretical paradigms in spatial prediction: UK as a benchmark stationary Gaussian process model, FRK as a reduced-rank approximation suitable for large datasets, and NNGP as a scalable local Gaussian process using conditional independence. The machine learning models complement geostatistical methods by relaxing distributional assumptions and capturing nonlinear relationships between predictors. This comparative design enables evaluation across the spectrum from fully parametric spatial models to nonparametric, data-driven algorithms.
3.2 Geostatistical Methods
Geostatistical methods are a set of statistical techniques used to analyze spatial/spatiotemporal data and explicitly take into account the spatial/spatiotemporal dependence of the data. These methods are widely used in various fields such as geology, hydrology, agriculture, environmental science, and many others (Maliva, 2016). A typical procedure in geostatistical techniques is to start with modeling the spatial/spatiotemporal correlation structure of the data, which describes how the values at different locations are related to each other in space and over time. Once we estimate the model parameters from the data, we can make predictions for the variable of interest at unsampled locations.
3.2.1 Universal Kriging
@@
where, $\mu(s,t)=\mbox{\bf X}(s,t)\mbox{$\beta$}$ is the mean with $\mbox{\bf X}(s,t)=[X_{1}(s,t),X_{2}(s,t),\dots,X_{p}(s,t)]^{T}$ are predictors and $\beta$ is the $p$-dimensional vector of parameters. $w(s,t)$ represents a mean-zero Gaussian process with a spatiotemporal covariance function. $\epsilon(s,t)$ is the white noise measurement error with mean 0 and variance $\sigma_{\epsilon}^{2}$. We use a separable exponential covariance function,
@@
where $\mbox{\bf h}=s^{\prime}-s$ is the spatial lag, $\tau=t^{\prime}-t$ temporal lag and $\mbox{$\theta$}=\{\sigma_{s},\sigma_{t},\rho_{s},\rho_{t}\}$ is the set of parameters associated with this covariance function, where $\sigma_{s},\sigma_{t}$ are the standard deviations for space and time , and $\rho_{s},\rho_{t}$ are spatial and time range parameters respectively. To predict Y at any unsampled location at $(s_{0},t_{0})$, we have
$\hat{Y}(s_{0},t_{0})=\mbox{\bf X}(s_{0},t_{0})^{T}\hat{\mbox{$\beta$}}_{gls}+c(\mbox{$\theta$})^{T}\Sigma_{y}^{-1}(\mbox{$\theta$})(\mbox{\bf Y}-\mbox{\bf X}\hat{\mbox{$\beta$}}_{gls}),$ (3)
where $c(\mbox{$\theta$})=\text{Cov}(Y(s_{0},t_{0}),\mbox{\bf Y})$ and $\text{Var}(\mbox{\bf Y})=\Sigma_{y}(\mbox{$\theta$})$. The uncertainty can be obtained by
$\text{Var}(\hat{Y}(s_{0},t_{0}))=c_{0,0}-c(\mbox{$\theta$})^{T}\Sigma_{y}^{-1}(\mbox{$\theta$})c(\mbox{$\theta$})+\kappa,$ (4)
@@
$\kappa=\left(\mbox{\bf X}(s_{0},t_{0})-\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{$\theta$})c(\mbox{$\theta$})\right)^{T}\left(\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{$\theta$})\mbox{\bf X}\right)^{-1}\left(\mbox{\bf X}(s_{0},t_{0})-\mbox{\bf X}^{T}\Sigma_{y}^{-1}(\mbox{$\theta$})c(\mbox{$\theta$})\right)$.
The above prediction method is called universal kriging. To obtain the prediction in Equation (3) and the prediction uncertainty in Equation (4), we compute the estimates $\hat{\mbox{$\beta$}}_{gls}$ and $\hat{\mbox{$\theta$}}$ to yield the empirical best linear unbiased predictor (EBLUP). Under the model specification, the random variable $Y(s_{0},t_{0})|\mathbf{Y}$ representing the value at an unobserved location conditional on the observed data, follows a Gaussian distribution. The mean of this distribution is given by Equation (3), $\hat{Y}(s_{0},t_{0})$, and its variance is given by Equation (4), $\text{Var}(\hat{Y}(s_{0},t_{0}))$. Consequently, a 95% prediction interval for $Y(s_{0},t_{0})$ is derived directly from this Gaussian predictive distribution and is given by:
$\hat{Y}(s_{0},t_{0})\pm Z_{0.975}\times\sqrt{\text{Var}(\hat{Y}(s_{0},t_{0}))},$ (5)
where $Z_{0.975}\approx 1.96$ is the 97.5th percentile of the standard normal distribution. In this study, we implemented universal kriging using the gstat package in R (Gräler et al., 2016)
3.2.2 Nearest Neighbor Gaussian Process